# Time-Correlated Video Bridge Matching

### v1

@@ 4.3 Choice of the Prior Process for Video Manipulation Tasks
Figure 3: Image-to-Video generation results. Based on the first frame, the methods need to generate the next 99 frames. As can be seen, TCVBM makes it possible to achieve better appearance of the generations and a lower quality degradation with increasing sequence length.
5 EXPERIMENTS
5.1 Implementation Details
To evaluate the performance of our new method and compare it to other approaches, we take our own small and simple U-Net model based on several residual blocks with $2D$ convolutions. The model has approximately $8.7$ million parameters. We conduct experiments on the MovingMNIST dataset (Srivastava et al., 2015) , which contains $10,000$ video sequences each $20$ frames long and showing $2$ moving digits in a $64\times 64$ resolution. For each experiment, we extract sub-sequences from the videos, consisting of $10$ consecutive frames, and concatenate them into a $10$-channel input for the neural network. We use $9,500$ training sequences and $500$ validation sequences, with a random split for each seed value. We train each model for $150,000$ iterations with a batch size of $128$ and an ema rate of $0.999$. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with betas set to $0.9$ and $0.95$, a weight decay of $10^{-4}$, and a learning rate of $3\times 10^{-5}$. The number of steps in the reverse process for all methods is equal to $1000$. All experiments are conducted using a single NVIDIA Tesla A100100 GPU.
@@ 5.1 Implementation Details
5.2 Experiments Roadmap
We compare TCVBM with three classical generative methods: DDPM (Ho et al., 2020), DDIM (Song et al., 2021a), and Bridge Matching with Brownian Bridge (BM) (Ibe, 2013, Chapter 9). For each method, we use the same setup, except for the loss function and sampling procedure in the forward and reverse processes, which are determined by the specific paradigm of each algorithm. Based on each of these four approaches, we train three models to solve the following tasks:
Frame Interpolation.
A sequence of $10$ frames is fed into the network as input, with the first and last frames from the dataset fixed. At the output, $8$ middle frames are generated, which are then compared with $8$ corresponding video frames from the dataset. The DDPM and DDIM methods aim to solve the generation problem for middle frames from random noise in the reverse process. For BM and TCVBM methods, we consider two ways to initialize the middle input frames: linear interpolation between the two fixed boundary frames, or noise samples from the Gaussian distribution $\mathcal{N}(\textbf{0},\mathbf{I})$. The second approach produced better results for BM and TCVBM, so the results presented here further are for noise-based initialization method. For more details, please see the Appendix C.1.
Image-to-Video Generation.
The task is to generate the remaining $9$ frames based on the first frame. We provide the first frame as input data, along with $9$ Gaussian samples from $\mathcal{N}(\textbf{0},\mathbf{I})$, or $9$ copies of the first frame. The second option demonstrated the best results for all methods, so further we consider this initialization variant. For more information, please see the Appendix C.2.
Video Super Resolution.
The network accepts $10$ low resolution frames, and $10$ frames with the dataset resolution are expected at the output. We consider starting with resolutions of $16\times 16$ or $32\times 32$ for all methods, but due to the simplicity of the task, we did not see any significant advantages for our method using the second option. Therefore, the further results are reported for the starting resolution of $16\times 16$. We also consider two initialization options: feeding the network with only low resolution data or low resolution data concatenated with Gaussian noise. Both approaches produced similar results, and here we use noise-free variant. More details can be found in the Appendix C.3.
Below, we present the results for the BM and TCVBM methods with $\epsilon$ equal to $0.1$ and a coefficient at $\mathbf{A}$ set to $1$ for the TCVBM model, i.e., $\widetilde{\mathbf{A}}=\alpha\mathbf{A}$ and $\widetilde{\mathbf{b}}=\alpha\mathbf{b}$, where $\alpha=1$. Additionally, we conduct a series of experiments to find optimal hyperparameters $\epsilon$ and $\alpha$. The results of these experiments can be found in Appendix D. Furthermore, we explored the dynamic correlation approach in which the coefficient $\alpha_{t}$ depends on time and increases as the reverse process progresses, i.e. when $t\rightarrow 0$. However, our experiments showed that this approach does not provide an advantage over TCVBM when the $\alpha$ is time-independent. The theoretical background and experimental findings are presented in Appendix E.
5.3 Results
Qualitative Comparison.
As can be seen in Figures 2, 3, and 4, TCVBM produces more consistent and accurate results compared to other generative methods. By considering the mutual correlation between frames, our approach ensures a smoother transition and information transfer between them, which enhances the overall method stability. Additional examples can be found in Appendix F.
Figure 4: Video super resolution results. The resolution is increased from $16\times 16$ to $64\times 64$. As can be seen, TCVBM outperforms other methods.
@@ 5.3 Results
We perform quantitative evaluation on $500$ videos from our MovingMNIST validation set using four metrics: FVD (Unterthiner et al., 2019), LPIPS (Zhang et al., 2018), PSNR, and SSIM. Tables 1, 2, and 3 present the metric values. As can be seen, in most cases TCVBM outperforms other approaches in terms of quantitative metrics. Additional results providing the standard deviation over launches with different random seeds are presented in the Appendix B.
@@ 6 DISCUSSION
This work has several limitations. First, the reliance on a predefined tridiagonal matrix prior is likely insufficient for capturing the complex nature of temporal dependencies in real-world data, particularly long-range and non-linear dynamics. Consequently, the exploration of more sophisticated interpolants – potentially modeling correlations in a feature space via adapters rather than directly between frames-is a critical direction for future work. Second, the computational cost of the iterative sampling process precludes real-time application. Finally, while the proposed approach offers a novel theoretical framework, its empirical validation is currently limited to the MovingMNIST dataset. Broader evaluation on large-scale, real-world datasets, complemented by human evaluation, is essential to substantiate its practical effectiveness and robustness.
Supplementary Materials
@@ Appendix A PROOF OF PROPOSITIONS
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$ (16)
@@
$\mathbf{I}_{t}:=\int_{0}^{t}e^{A(t-s)}\,d\mathbf{W}_{s},\qquad\text{so that}\qquad\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=\sqrt{\epsilon}\,\mathbf{I}_{t}.$
$\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\mathbb{E}\big[(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0})(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0})^{\top}\mid\mathbf{X}_{0}\big]=\epsilon\,\mathrm{Cov}(\mathbf{I}_{t}).$
@@
$\boldsymbol{\Sigma}_{t|0}=\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1}.$
@@
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}\mid\mathbf{X}_{0})=-\,\boldsymbol{\Sigma}_{t|0}^{-1}\big(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\big).$
@@
Step 1: Joint law from the prior. From (15) and (16) in the proof of Proposition 1:
$\mathbf{X}_{u}=\boldsymbol{\mu}_{u|0}(\mathbf{X}_{0})+\sqrt{\epsilon}\int_{0}^{u}e^{\mathbf{A}(u-s)}d\mathbf{W}_{s},$
$\boldsymbol{\mu}_{u|0}(\mathbf{X}_{0})=e^{\mathbf{A}u}\mathbf{X}_{0}+\big(e^{\mathbf{A}u}-\mathbf{I}\big)\mathbf{A}^{-1}\mathbf{b}.$
@@
$\mathbb{E}[\mathbf{X}_{t}\mid\mathbf{X}_{0}]=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}),\qquad\mathbb{E}[\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0}]=\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0}),$
$\boldsymbol{\Sigma}_{t|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1},\qquad\boldsymbol{\Sigma}_{t^{\prime}|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t^{\prime}}-\mathbf{I}\big)\mathbf{A}^{-1}.$
@@
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}=\mathrm{Cov}(\mathbf{X}_{t},\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0})=\epsilon\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}(t^{\prime}-s)}\,ds=\epsilon\int_{0}^{t}e^{\mathbf{A}(t+t^{\prime}-2s)}\,ds=\frac{\epsilon}{2}\,\mathbf{A}^{-1}\!\left(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\right).$
@@
$\boldsymbol{\mu}_{t\mid 0,t^{\prime}}=\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})+\boldsymbol{\Sigma}_{t,t^{\prime}|0}\,\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\!\big(\mathbf{X}_{t^{\prime}}-\boldsymbol{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})\big),$
$\boldsymbol{\Sigma}_{t\mid 0,t^{\prime}}=\boldsymbol{\Sigma}_{t|0}-\boldsymbol{\Sigma}_{t,t^{\prime}|0}\,\boldsymbol{\Sigma}_{t^{\prime}|0}^{-1}\boldsymbol{\Sigma}_{t,t^{\prime}|0}.$
@@
$\boldsymbol{\Sigma}_{t|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t}-\mathbf{I})\mathbf{A}^{-1},\quad\boldsymbol{\Sigma}_{t^{\prime}|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t^{\prime}}-\mathbf{I})\mathbf{A}^{-1},\quad\boldsymbol{\Sigma}_{t,t^{\prime}|0}=\tfrac{\epsilon}{2}A^{-1}\big(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\big).$
@@
$v_{\phi}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)$
@@
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)+\boldsymbol{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\boldsymbol{\Sigma}_{t|0}^{-1}\left((\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0}))-\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)^{\top}(\boldsymbol{\Sigma}_{t|0}^{\top})^{-1}\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]$
@@
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[2(\boldsymbol{\Sigma}_{t|0}^{\top})^{-1}\boldsymbol{\Sigma}_{t|0}^{-1}\left(\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]=0$
Since $\boldsymbol{\Sigma}_{t|0}^{-1}$ is positive definite we can multiply by $\frac{1}{2}\boldsymbol{\Sigma}_{t|0}(\boldsymbol{\Sigma}_{t|0}^{\top})$ and get:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
@@
$\mathbb{E}_{\mathbf{X}_{t},t}\left[\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
From (16) we have:
@@
$v^{*}(\mathbf{X}_{t},t)=-\boldsymbol{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right)$
@@
Appendix B ADDITIONAL QUANTITATIVE RESULTS

### v2

@@ 5.2 Results
We perform quantitative evaluation on $500$ videos from our MovingMNIST validation set using FVD (Unterthiner et al., 2019), LPIPS (Zhang et al., 2018), PSNR, and SSIM. As shown in Tables 1, 2, and 3, TCVBM outperforms other methods on most metrics. Additional results, including standard deviations across random seeds, are in Appendix C.
@@ 6 Discussion
This work has several limitations. First, the predefined tridiagonal prior may be insufficient for complex, long-range, or non-linear temporal dependencies in real-world data; future work should explore more sophisticated interpolants, such as those modeling correlations in feature space. Second, due to the computational cost, empirical validation is currently limited to MovingMNIST dataset. Broader evaluation on large-scale real-world datasets and human assessment are needed to confirm its practical effectiveness and robustness.
@@ Appendix A Proof of Propositions
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\Big(\!\int_{0}^{t}e^{\mathbf{A}(t-s)}ds\Big)\mathbf{b}=e^{\mathbf{A}t}X_{0}+\big(e^{\mathbf{A}t}-I\big)\mathbf{A}^{-1}\mathbf{b},$ (16)
@@
$\mathbf{I}_{t}:=\int_{0}^{t}e^{A(t-s)}\,d\mathbf{W}_{s},\qquad\text{so that}\qquad\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})=\sqrt{\epsilon}\,\mathbf{I}_{t}.$
$\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\mathbb{E}\big[(\mathbf{X}_{t}-\bm{\mu}_{t|0})(\mathbf{X}_{t}-\bm{\mu}_{t|0})^{\top}\mid\mathbf{X}_{0}\big]=\epsilon\,\mathrm{Cov}(\mathbf{I}_{t}).$
@@
$\bm{\Sigma}_{t|0}=\mathrm{Cov}(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1}.$
@@
$\nabla_{\mathbf{X}_{t}}\log q(\mathbf{X}_{t}\mid\mathbf{X}_{0})=-\,\bm{\Sigma}_{t|0}^{-1}\big(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0})\big).$
@@
Step 1: Joint law from the prior. From equation 15 and equation 16 in the proof of Proposition 1:
$\mathbf{X}_{u}=\bm{\mu}_{u|0}(\mathbf{X}_{0})+\sqrt{\epsilon}\int_{0}^{u}e^{\mathbf{A}(u-s)}d\mathbf{W}_{s},$
$\bm{\mu}_{u|0}(\mathbf{X}_{0})=e^{\mathbf{A}u}\mathbf{X}_{0}+\big(e^{\mathbf{A}u}-\mathbf{I}\big)\mathbf{A}^{-1}\mathbf{b}.$
@@
$\mathbb{E}[\mathbf{X}_{t}\mid\mathbf{X}_{0}]=\bm{\mu}_{t|0}(\mathbf{X}_{0}),\qquad\mathbb{E}[\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0}]=\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0}),$
$\bm{\Sigma}_{t|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t}-\mathbf{I}\big)\mathbf{A}^{-1},\qquad\bm{\Sigma}_{t^{\prime}|0}=\frac{\epsilon}{2}\big(e^{2\mathbf{A}t^{\prime}}-\mathbf{I}\big)\mathbf{A}^{-1}.$
@@
$\bm{\Sigma}_{t,t^{\prime}|0}=\mathrm{Cov}(\mathbf{X}_{t},\mathbf{X}_{t^{\prime}}\mid\mathbf{X}_{0})=\epsilon\int_{0}^{t}e^{\mathbf{A}(t-s)}\,e^{\mathbf{A}(t^{\prime}-s)}\,ds=\epsilon\int_{0}^{t}e^{\mathbf{A}(t+t^{\prime}-2s)}\,ds=\frac{\epsilon}{2}\,\mathbf{A}^{-1}\!\left(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\right).$
@@
$\bm{\mu}_{t\mid 0,t^{\prime}}=\bm{\mu}_{t|0}(\mathbf{X}_{0})+\bm{\Sigma}_{t,t^{\prime}|0}\,\bm{\Sigma}_{t^{\prime}|0}^{-1}\!\big(\mathbf{X}_{t^{\prime}}-\bm{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})\big),$
$\bm{\Sigma}_{t\mid 0,t^{\prime}}=\bm{\Sigma}_{t|0}-\bm{\Sigma}_{t,t^{\prime}|0}\,\bm{\Sigma}_{t^{\prime}|0}^{-1}\bm{\Sigma}_{t,t^{\prime}|0}.$
@@
$\bm{\Sigma}_{t|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t}-\mathbf{I})\mathbf{A}^{-1},\quad\bm{\Sigma}_{t^{\prime}|0}=\tfrac{\epsilon}{2}(e^{2\mathbf{A}t^{\prime}}-\mathbf{I})\mathbf{A}^{-1},\quad\bm{\Sigma}_{t,t^{\prime}|0}=\tfrac{\epsilon}{2}A^{-1}\big(e^{\mathbf{A}(t+t^{\prime})}-e^{\mathbf{A}(t^{\prime}-t)}\big).$
@@
$v_{\phi}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)$
@@
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|v_{\phi}(\mathbf{X}_{t},t)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right],$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)+\bm{\Sigma}_{t|0}^{-1}(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\bm{\Sigma}_{t|0}^{-1}\left((\mathbf{X}_{t}-\bm{\mu}_{t|0}(\mathbf{X}_{0}))-\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))\right)\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left\|\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right\|^{2}\right]=$
$\min_{\phi}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)^{\top}(\bm{\Sigma}_{t|0}^{\top})^{-1}\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]$
@@
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[2(\bm{\Sigma}_{t|0}^{\top})^{-1}\bm{\Sigma}_{t|0}^{-1}\left(\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right)\right]=0$
Since $\bm{\Sigma}_{t|0}^{-1}$ is positive definite we can multiply by $\frac{1}{2}\bm{\Sigma}_{t|0}(\bm{\Sigma}_{t|0}^{\top})$ and get:
$\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{t},t}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
@@
$\mathbb{E}_{\mathbf{X}_{t},t}\left[\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]\right]=0$
$\mathbb{E}_{\mathbf{X}_{0}|\mathbf{X}_{t}}\left[\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t},t))-\bm{\mu}_{t|0}(\mathbf{X}_{0})\right]=0$
From equation 16 we have:
@@
$v^{*}(\mathbf{X}_{t},t)=-\bm{\Sigma}_{t|0}^{-1}\left(\mathbf{X}_{t}-\bm{\mu}_{t|0}(\widehat{\mathbf{X}}_{0}^{*}(\mathbf{X}_{t},t))\right)$
@@
Appendix B Implementation Details
To evaluate the performance of our method and compare it to other approaches, we take our own small and simple U-Net model based on several residual blocks with $2D$ convolutions. The model has approximately $8.7$ million parameters. For each experiment, we extract sub-sequences from the videos, consisting of $10$ consecutive frames, and concatenate them into a $10$-channel input for the neural network. We use $9,500$ training sequences and $500$ validation sequences, with a random split for each seed value. We train each model for $150,000$ iterations with a batch size of $128$ and an ema rate of $0.999$. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with betas set to $0.9$ and $0.95$, a weight decay of $10^{-4}$, and a learning rate of $3\times 10^{-5}$. The number of steps in the reverse process for all methods is equal to $1000$. All experiments are conducted using a single NVIDIA Tesla A100100 GPU.
Appendix C Additional Quantitative Results